If 0.5 mol of BaCl2 is mixed with 0.20 mol of Na3PO4, the maximum amount of Ba3(PO4)2 that can be formed is


(a) 0.70 mol (b) 0.50 mol (c) 0.20 mol (d) 0.1 mol


please explain how?

Grade:12

4 Answers

Vikas TU
10 years ago

I think it''s a mole -1 chptr. qstn.

 

             3BaCl2 + 2Na3PO4  ------------->  Ba3(PO4)2 + 6Nacl

                  0.5 mol           0.20 mol               x mol           

                First find which is the limiting reagent?

              0.5/3 = 1/6          0.2/2 = 1/10

                            1/10 < 1/6

              thus,  Na3PO4 is the limiting reagent.

               Now, 1/10 = x/1

                             x = 0.1 mol.

 

(d) may be the ansr.

plz approvE!

Roshan Mohanty
10 years ago

3 BaCl2 + 2 Na3PO4 = Ba3(PO4)2 + 6NaCl
Here clearly BaCl2 is the limiting agent
So 3 moles of bacl2 gives 1 mole of ba3(po4)2
so 0.5 mole will give 0.2 mole(approx)

I hope i am correct 

Saksham Soni
6 years ago
3bacl2 +2na3po4=ba3(po4)2+6naclHear clearly NA3PO4 is limiting reagentNow 1/10=x/1 X=0.1molSo D is the answer
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

3 BaCl2 + 2 Na3PO4 -------------> Ba3(PO4)2 + 6 NaCl
0.5 mol 0.20 mol x mol
Lets first find the limiting reagent?
0.5/3 = 1/6 0.2/2 = 1/10
1/10 < 1/6
thus, Na3PO4 is the limiting reagent.
Now, 1/10 = x/1
x = 0.1 mol.

Thanks and Regards